In the spring of​ 2016, the CFO of HTPL Distributing Company decided to distribute a stock dividend to its shareholders.​ Specifically, the CFO proposed that the company pay 0.055 shares of stock to the holders of each share of common stock such that the holder of​ 1,000 shares of stock would receive an additional 55 shares of common stock.

a. If the firm had total net income for the year of $10,000,000 and had 22,000,000 shares of common stock outstanding before the stock​ dividend, what was the​ firm's earnings per​ share? b. After paying the stock​ dividend, what was the​ firm's earnings per​ share?

c. If you owned​ 1,000 shares of stock before the stock​ dividend, how many dollars of earnings did the firm earn from your​ 1,000-share investment? After the stock dividend was​ paid, how many dollars of earnings did the firm earn on your larger share​ holdings? What effect would you expect from the payment of the stock dividend on your total investment in the​ firm?
Answer #1

a). EPS = net income/shares outstanding

Before stock dividend, EPS = 10,000,000/22,000,000 = 0.4545 per share

b). After stock dividend, number of shares becomes (1,055/1,000)*22,000,000 = 23,210,000

EPS = 10,000,000/23,210,000 = 0.4308 per share

c). Before stock dividend, earnings on your investment = number of shares*EPS = 1,000*0.4545 = $454.55

After stock dividend, earnings on your investment = 1,055*0.4308 = $454.55